# MoEMeta: Mixture-of-Experts Meta Learning for Few-Shot Relational Learning

### v1

@@ 4.3 Task-Tailored Local Adaptation
Specifically, for each task, MoEMeta maintains a task-specific set of local parameters, denoted as $\bm{\eta}$, including three randomly initialized projection vectors, $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$, which project the embeddings $\mathbf{h}_{j}$, $\mathbf{R}_{\mathcal{T}_{r}}$, and $\mathbf{t}_{j}$ into a task-specific space. Inspired by TransD (Ji et al., 2015), which addresses the limitations of fixed mapping strategies using per-entity and per-relation projections, our approach leverages task-specific projections to avoid overparametrization under few-shot settings. As all triplets within a task share the same relation, we adopt a unified projection vector $\mathbf{p}_{h}$ for all head entities, and $\mathbf{p}_{t}$ for all tail entities within the task. This provides a lightweight solution for enabling efficient task-specific adaptation with minimal parameter overhead.
@@
Above, the dot product $\mathbf{p}^{\top}\mathbf{R}_{\mathcal{T}_{r}}$ yields a scalar alignment score between each projection vector $\mathbf{p}$ and relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which then scales $\mathbf{R}_{\mathcal{T}_{r}}$ to generate an offset vector that modulates specific features in the entity embeddings and relation-meta. Although the three projection vectors can functionally be merged into a single perturbation term, maintaining separate projections for head entities, relation, and tail entities enhances modularity and flexibility, allowing for more expressive or heterogeneous projection functions that can incorporate richer inductive biases or task-specific modeling needs in future extensions. The updated embeddings ($\mathbf{h}_{j}^{\prime}$, $\mathbf{t}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$) thus capture task-specific nuances while maintaining consistency with the generalizable meta-knowledge. This local adaptation scheme ensures that each task can adaptively focus on its unique aspects but also preserve the shared commonalities across related tasks.
@@
where $\lVert\cdot\rVert_{2}$ denotes the $\ell_{2}$ norm. $\mathbf{h}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{j}^{\prime}$ are the projected embeddings of the head entity, relation, and tail entity, respectively. This formulation ensures that the distance between the adapted head entity embedding $\mathbf{h}_{j}^{\prime}$ and the adapted tail entity embedding $\mathbf{t}_{j}^{\prime}$ is minimized after being translated by the relation embedding $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$ in the task-specific projection space.
Considering the entire support set, we define a margin-based loss function as follows:
@@
Once the gradients of the loss function $\mathcal{L}(\mathcal{S}_{r})$ are computed, we optimize the projection vectors $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, $\mathbf{p}_{t}$, together with the relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, to ensure generalization to the query set. Specifically, the projection vectors can be updated in a unified manner as follows:
@@
where $\alpha$ is the learning rate. These updates adapt the model based on the support set, which is then evaluated on the query set $\mathcal{Q}_{r}$. For each query triplet $(h_{q},r,t_{q})$, we evaluate its plausibility for each $t_{q}$ from the candidate set $C_{h_{q},r}$, using the score function Eq. (12). Then, the query loss is defined as:
@@
The task-specific projection vectors $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$ are randomly initialized and trained specifically for each task on the support set $\mathcal{S}_{r}$, enabling the model to adapt locally to each task. This mechanism ensures that (1) task-specific adaptation: By projecting embeddings into a task-specific latent space, the model captures fine-grained task-specific local contexts. (2) complementarity to MoE: While the MoE captures globally shared relational prototypes, dynamic learnable projections provide finer-grained local adaptation, thereby improving model generalization under few-shot settings.
@@ 4.4 MoEMeta Meta-Training and Meta-Testing
During meta-testing, the learned global parameters $\bm{\Phi}$ are used to initialize the model and remain frozen. For each novel task, only the local parameters $\bm{\eta}$ are fine-tuned on the support set before evaluation on the query set by ranking candidate tail entities, following the same inference procedure as in meta-training. The complexity analysis of MoEMeta’s components are provided in Appendix A.
Input: Knowledge Graph $\mathcal{G}$
Output: Optimized model parameters $\bm{\Phi}$
@@
2Initialize globally optimized parameters $\bm{\Phi}=\{\textbf{W},\bm{\beta},\{\bm{\theta}\}_{i=1}^{M},\bm{\theta}_{g}\}$;
@@
5   Sample $\mathcal{T}_{r}$ = $\mathcal{S}_{r}$, $\mathcal{Q}_{r}$};
6 foreach triplet $(h_{i},r,t_{i})\in\mathcal{S}_{r}$ do
@@ 5   Sample $\mathcal{T}_{r}$ = $\mathcal{S}_{r}$, $\mathcal{Q}_{r}$};
7      Attentively aggregate neighbors using Eq. 2 -Eq. 4;
@@
8      Compute expert selection scores using Eq. 7;
9      Select top-NN experts and compute $\mathbf{r}_{i}$;
@@
11   end foreach
12  Compute relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ using Eq. 8;
13
 // Task-Tailored Local Adaptation
14   Initialize locally optimized parameters $\bm{\eta}=\{\mathbf{p}_{h},\mathbf{p}_{r},\mathbf{p}_{t}\}$;
15 foreach triplet $(h_{i},r,t_{i})\in\mathcal{S}_{r}$ do
@@ 14   Initialize locally optimized parameters $\bm{\eta}=\{\mathbf{p}_{h},\mathbf{p}_{r},\mathbf{p}_{t}\}$;
17
18   end foreach
 // Inner loop optimization
19   Compute the support loss $\mathcal{L}(\mathcal{S}_{r})$ using Eq. 13;
20   Optimize projection vectors using Eq. 14;
21   Optimize relation-meta representation using Eq. 15;
22
 // Evaluation on Query Set
23   Compute the query loss for all $(h_{q},r,t_{q})\in\mathcal{Q}_{r}$ via Eq. 16;
 // Optimization
24   Optimize model parameters $\bm{\Phi}$ based on the query loss.
25 end while
@@ 5 Experiments
Datasets. We evaluate our method on three widely used KG benchmarks for few-shot relational learning: Nell-One and Wiki-One (Xiong et al., 2018), as well as FB15K-One (Ran et al., 2024). The training/validation/test splits include 51/5/11, 133/16/34, and 75/11/33 tasks on Nell-One, Wiki-One, and FB15K-One, respectively. Detailed dataset statistics are summarized in Table 1.
@@
Evaluation Metrics. We report both MRR (mean reciprocal rank) and Hits@k ($k=1,5,10$). MRR captures the average reciprocal rank of correct entities, while Hits@k measures the proportion of correct entities appearing in the top $k$ predictions. Following prior studies, our method is evaluated against baselines under 1-shot and 5-shot settings on Nell-One and Wiki-One, which are commonly used for evaluating FSRL, and under the 3-shot setting on FB15K-One, following the setup of RelAdapter (Ran et al., 2024).
@@
  • Conventional KG embedding methods: This group focuses on modeling relational structures in KGs to learn entity and relation embeddings. We compare with the widely-used baselines including TransE (Bordes et al., 2013), TransD (Ji et al., 2015), DistMult (Yang et al., 2015), and ComplEx (Trouillon et al., 2016). Using OpenKE (Han et al., 2018), these models are reproduced with hyperparameters specified in the original papers.
  • Few-shot relational learning methods: We compare our method against state-of-the-art FSRL baselines evaluated under the same framework, (1) metric learning methods: GMatching (Xiong et al., 2018), FAAN (Sheng et al., 2020), FSKGC (Zhang et al., 2020); and (2) strong meta-learning competitors: MetaR (Chen et al., 2019), GANA (Niu et al., 2021), HiRe (Wu et al., 2023), and RelAdapter (Ran et al., 2024). We exclude methods like CSR (Huang et al., 2022) and NP-FKGC (Luo et al., 2023) as they employ a different training and evaluation protocol.
All results are obtained after the models are trained using all triplets from background relations (Xiong et al., 2018) and meta-training tasks, and evaluated on meta-testing tasks.
@@ 5.1 Comparison with Baseline Methods
Tables 2 and 3 compare MoEMeta against state-of-the-art methods on Nell-One and Wiki-One under 1-shot and 5-shot settings, and on FB15K-One under 3-shot setting. The best and the second best performers are respectively marked by grey and underline. Our proposed MoEMeta consistently achieves state-of-the-art results across all datasets and settings. On Nell-One, MoEMeta surpasses HiRe, the second best performer, by 11.8% MRR, 23.9% Hits@1, 6.4% Hits@5, and 6.1% Hits@10 under $1$-shot setting, and 10.8% MRR, 14.0% Hits@1, 2.3% Hits@5, and 0.6% Hits@10 under $5$-shot setting. On FB15K-One, MoEMeta outperforms the second best method by 4.4% MRR, 0.3% Hits@1, and 13.2% Hits@5 under $3$-shot setting. We further conduct paired t-tests between MoEMeta and the second best performers on all datasets. The improvements of MoEMeta over the second best method with statistical significance at the 0.05 level are marked with . These results highlight MoEMeta’s effectiveness in learning generalizable meta-knowledge that adapts well to novel relations and proving its competitiveness against other baselines.
@@ 5.2 Ablation Studies
To assess the contribution of MoEMeta’s three key components, attentive neighbor aggregation, mixture of experts (MoE), and task-tailored local adaptation, we report the ablation study of MoEMeta on Nell-One under 1-shot and 5-shot settings, and FB15K-One under 3-shot setting in Table 4.
  • Removing Neighbor Aggregation (w/o N.A), which removes the first term in Eq. 4, results in a moderate performance decrease across all settings. This suggests the importance of incorporating neighboring information to enhance entity representations.
@@ 5.3 Analysis on Complex Relations
To further evaluate the significance of task-tailored local adaptation for complex relations, we conduct a detailed analysis on Nell-One under the 5-shot setting by categorizing meta-testing set into three distinct relation types: 1-N, N-1, and N-N. We compare MoEMeta with a variant without task-tailored local adaptation (w/o L.A). As shown in Table 4, the removal of local adaptation leads to a substantially larger performance drop on 1-N and N-N relations than on N-1 relations. This is because the tail entity in N-1 relations is relatively predictable, reducing the search space and enabling the model to learn a more stable mapping pattern. In contrast, 1-N and N-N relations exhibit greater ambiguity and a more dispersed distribution of tail entities, making them harder to predict. Notably, removing task-tailored adaptation results in the most significant performance drop in the N-N relation type. This analysis confirms the critical role of our task-tailored adaptation mechanism in enhancing generalization to complex relational patterns.
For better elaboration, we further visualize the distribution of gating values of different relations in Figure 2 to provide a more intuitive understanding of why learning relational prototypes would work. As observed, relations with similar patterns exhibit a comparable gating value distribution and activate a similar subset of experts, whereas distinct relations demonstrate an opposite pattern, highlighting the model’s ability to differentiate between diverse relational patterns but capture similar relational prototypes that are globally shared across related tasks.
@@ 5.4 Hyperparameter Analysis on MoE
To assess their impact of two MoE hyperparameters, the number of experts $M$ and the number of top selected experts $N$, we conduct a sensitivity analysis on Nell-One under 5-shot settings as a case study. When fixing $N=5$ and varying $M\in\{16,32,64\}$, MoEMeta achieves the best performance with 32 experts. This aligns with the guideline of setting the number of experts $M$ smaller than the total number of tasks to encourage learning shared relational prototypes and prevent overfitting with limited support triplets. When fixing $M=32$ and varying $N\in\{3,5,10\}$, we observe that selecting slightly more experts (e.g., 5 vs. 3) leads to improved performance, yet increasing $N$ to 10 causes only a slight drop, suggesting that MoEMeta is relatively insensitive to $N$ within a reasonable range.

### v2

@@ 4.3 Task-Tailored Local Adaptation
Specifically, for each task, MoEMeta maintains a task-specific set of local parameters, denoted as $\bm{\eta}$, namely three randomly initialized projection vectors, $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$, which project the embeddings $\mathbf{h}_{j}$, $\mathbf{R}_{\mathcal{T}_{r}}$, and $\mathbf{t}_{j}$ into a task-specific subspace. Inspired by TransD (Ji et al., 2015), which addresses the limitations of fixed mapping strategies using per-entity and per-relation projections, our approach leverages task-specific projections to avoid overparametrization under few-shot settings. As all triplets within a task share the same relation, we adopt a unified projection vector $\mathbf{p}_{h}$ for all head entities, and $\mathbf{p}_{t}$ for all tail entities within the task. This provides a lightweight solution for enabling efficient task-specific adaptation with minimal parameter overhead.
@@
Above, the dot product $\mathbf{p}^{\top}\mathbf{R}_{\mathcal{T}_{r}}$ yields a scalar alignment score between each projection vector $\mathbf{p}$ and relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which then scales $\mathbf{R}_{\mathcal{T}_{r}}$ to generate an offset vector that modulates specific features in the entity embeddings and relation-meta. The updated embeddings ($\mathbf{h}_{j}^{\prime}$, $\mathbf{t}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$) thus capture task-specific nuances while maintaining consistency with the generalizable meta-knowledge. Although the three projection vectors could functionally be merged into a single perturbation term, maintaining separate projections for the head entity, relation-meta, and tail entity enhances modularity and flexibility. This design allows for more expressive or heterogeneous projection functions that can incorporate richer inductive biases or task-specific modeling needs in future extensions. Overall, this local adaptation scheme ensures that each task can adaptively focus on its unique aspects while preserving shared commonalities across related tasks.
@@
where $\lVert\cdot\rVert_{2}$ denotes the $\ell_{2}$ norm. $\mathbf{h}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{j}^{\prime}$ are the projected embeddings of the head entity, relation-meta, and tail entity, respectively. This formulation ensures that, in the task-specific projection space, the distance between the adapted head entity embedding $\mathbf{h}_{j}^{\prime}$ and the adapted tail entity embedding $\mathbf{t}_{j}^{\prime}$ is minimized after being translated by the relation-meta embedding $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$.
Consider the entire support set, we define a margin-based loss function as follows:
@@
Once the gradients of the support loss $\mathcal{L}(\mathcal{S}_{r})$ are computed, we update the projection vectors $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, $\mathbf{p}_{t}$, together with the relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, to ensure generalization to the support set. Specifically, the projection vectors can be updated in a unified manner as follows:
@@
where $\alpha$ is the learning rate. These updates adapt the model using the support set, which is then evaluated on the query set $\mathcal{Q}_{r}$. For each query triplet $(h_{q},r,t_{q})\in\mathcal{Q}_{r}$, we evaluate its plausibility for each $t_{q}$ from the candidate set $C_{h_{q},r}$, using the score function Eq. (12). Then, the query loss is similarly defined as:
@@
The task-specific projection vectors $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$ are randomly initialized and trained for each task on the support set $\mathcal{S}_{r}$, enabling effective local adaptation. This mechanism provides two key advantages. (1) Task-specific adaptation: By projecting embeddings into a task-specific latent space, the model captures fine-grained task-specific local contexts. (2) Complementarity to MoE: While the MoE captures globally shared relational prototypes, dynamic learnable projections provide finer-grained local adaptation, thereby improving model generalization under few-shot settings.
@@ 4.4 MoEMeta Meta-Training and Meta-Testing
During meta-testing, the learned global parameters $\bm{\Phi}$ are used to initialize the model and remain frozen. For each novel task, only the local parameters $\bm{\eta}$ are fine-tuned on the support set before evaluation on the query set by ranking candidate tail entities, following the same inference procedure as in meta-training. The complexity analysis of MoEMeta’s components is provided in Appendix A.
Input: Meta-training task set $\mathcal{D}_{\text{tr}}$, knowledge graph $\mathcal{G}$
Output: Optimized global parameters $\bm{\Phi}$
@@
2Initialize global parameters $\bm{\Phi}=\{\textbf{W},\bm{\beta},\{\bm{\theta}\}_{i=1}^{M},\bm{\theta}_{g}\}$;
@@
5   Sample a task $\mathcal{T}_{r}=\{\mathcal{S}_{r},\mathcal{Q}_{r}\}\in\mathcal{D}_{\text{tr}}$;
6 foreach $(h_{i},r,t_{i})\in\mathcal{S}_{r}$ do
@@ 5   Sample a task $\mathcal{T}_{r}=\{\mathcal{S}_{r},\mathcal{Q}_{r}\}\in\mathcal{D}_{\text{tr}}$;
7      Attentively aggregate neighbors using Eq. 2–Eq. 4;
@@
8      Compute expert scores via Eq. 7;
9      Select top-NN experts and compute relation representation $\mathbf{r}_{i}$;
@@
11  Compute the initial relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ via Eq. 8;
12
 // Inner Loop: Task-Tailored Local Adaptation
13   Initialize task-specific adaptation parameters $\bm{\eta}=\{\mathbf{p}_{h},\mathbf{p}_{r},\mathbf{p}_{t}\}$;
14 foreach $(h_{i},r,t_{i})\in\mathcal{S}_{r}$ do
@@
16
17  Compute the support loss $\mathcal{L}(\mathcal{S}_{r})$ via Eq. 13;
18   Update task-specific projection vectors $\mathbf{p}_{h},\mathbf{p}_{r},\mathbf{p}_{t}$ via Eq. 14;
19   Update relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ via Eq. 15;
 // Outer Loop: Meta-Update
20   Compute the query loss $\mathcal{L}(\mathcal{Q}_{r})$ over all $(h_{q},r,t_{q})\in\mathcal{Q}_{r}$ via Eq. 16;
21   Update global parameters $\bm{\Phi}$ via gradient descent on $\mathcal{L}(\mathcal{Q}_{r})$.
@@ 5 Experiments
Datasets. We evaluate our method on three widely used KG benchmarks specifically for few-shot relational learning: Nell-One and Wiki-One (Xiong et al., 2018), as well as FB15K-One (Ran et al., 2024). The training/validation/test splits include 51/5/11, 133/16/34, and 75/11/33 tasks on Nell-One, Wiki-One, and FB15K-One, respectively. Detailed dataset statistics are summarized in Table 1.
@@
Evaluation Metrics. We report both MRR (mean reciprocal rank) and Hits@$k$ ($k=1,5,10$). MRR captures the average reciprocal rank of correct entities, while Hits@k measures the proportion of correct entities appearing in the top $k$ predictions. Following prior studies, our method is evaluated against baselines under the 1-shot and 5-shot settings on Nell-One and Wiki-One, which are commonly used for evaluating FSRL, and under the 3-shot setting on FB15K-One, following the setup of RelAdapter (Ran et al., 2024).
@@
  • Conventional KG embedding methods: This group focuses on modeling relational structures in KGs to learn entity and relation embeddings. We compare with four widely-used baselines including TransE (Bordes et al., 2013), TransD (Ji et al., 2015), DistMult (Yang et al., 2015), and ComplEx (Trouillon et al., 2016). Using OpenKE (Han et al., 2018), these models are reproduced with hyperparameters specified in the original papers.
  • Few-shot relational learning methods: We compare our method against state-of-the-art FSRL baselines evaluated under the same framework, including (1) metric-learning methods: GMatching (Xiong et al., 2018), FAAN (Sheng et al., 2020), FSKGC (Zhang et al., 2020); and (2) strong meta-learning competitors: MetaR (Chen et al., 2019), GANA (Niu et al., 2021), HiRe (Wu et al., 2023), and RelAdapter (Ran et al., 2024). We exclude methods like CSR (Huang et al., 2022) and NP-FKGC (Luo et al., 2023) as they employ a different training and evaluation protocol.
All results are obtained after models are trained using all triplets from background relations (Xiong et al., 2018) and meta-training tasks, and then evaluated on meta-testing tasks.
@@ 5.1 Comparison with Baseline Methods
Tables 2 and 3 compare MoEMeta against state-of-the-art methods on Nell-One and Wiki-One under the 1-shot and 5-shot settings, and on FB15K-One under the 3-shot setting. The best and the second best performers are respectively marked by grey and underline. Our proposed MoEMeta consistently achieves state-of-the-art results across all datasets and settings. On Nell-One, MoEMeta surpasses HiRe, the second best performer, by 11.8% MRR, 23.9% Hits@1, 6.4% Hits@5, and 6.1% Hits@10 under $1$-shot setting, and 10.8% MRR, 14.0% Hits@1, 2.3% Hits@5, and 0.6% Hits@10 under $5$-shot setting. On FB15K-One, MoEMeta outperforms the second best method by 4.4% MRR, 0.3% Hits@1, and 13.2% Hits@5 under $3$-shot setting. We further conduct paired t-tests between MoEMeta and the second best performers on all datasets. The improvements of MoEMeta over the second best method with statistical significance at the 0.05 level are marked with . These results highlight MoEMeta’s effectiveness in learning generalizable meta-knowledge that adapts well to novel relations and proving its competitiveness against other baselines.
@@ 5.2 Ablation Studies
To assess the contribution of MoEMeta’s three key components, attentive neighbor aggregation, MoE-based meta-knowledge learning, and task-tailored local adaptation, we report ablation results on Nell-One under the 1-shot and 5-shot settings, and FB15K-One under the 3-shot setting in Table 4.
  • Removing attentive Neighbor Aggregation (w/o N.A), which removes the first term in Eq. 4, results in a moderate performance decrease across all settings. This suggests the importance of incorporating neighboring information to enhance entity representations.
@@ 5.3 Analysis on Complex Relations
To further evaluate the significance of task-tailored local adaptation for complex relations, we conduct a detailed analysis on Nell-One under the 5-shot setting by categorizing the meta-testing set into three relation types: 1-N, N-1, and N-N. In particular, we compare MoEMeta with a variant that removes task-tailored local adaptation (w/o L.A). As shown in Table 4, the removal of local adaptation leads to a substantially larger performance drop on 1-N and N-N relations than on N-1 relations. This is because, in N-1 relations, the tail entity is relatively predictable, with a smaller search space and more stable mapping patterns. In contrast, 1-N and N-N relations exhibit higher ambiguity and a more dispersed distribution of tail entities, making them more difficult to predict. Notably, removing task-tailored adaptation results in the largest performance drop for the N-N relation type. This analysis confirms the critical role of our task-tailored adaptation mechanism in enhancing generalization to complex and highly ambiguous relational patterns.
In addition, we visualize the distributions of gating values of different relations in Figure 2 to provide a more intuitive understanding of why learning relational prototypes works. As observed, relations with similar patterns (in the upper figure) exhibit a comparable gating value distribution and activate similar subsets of experts, whereas dissimilar relations (in the lower figure) demonstrate a markedly different pattern. This observation highlights our model’s ability to distinguish diverse relational patterns but capture globally shared relational prototypes across related tasks.
@@ 5.4 Hyperparameter Analysis on MoE
To assess the impact of two MoE hyperparameters, the number of experts $M$ and the number of top selected experts $N$, we conduct a sensitivity analysis on Nell-One under the 5-shot setting as a case study. When fixing $N=5$ and varying $M\in\{16,32,64\}$, MoEMeta achieves the best performance with 32 experts. This aligns with the guideline of setting the number of experts $M$ smaller than the total number of tasks to encourage learning shared relational prototypes and prevent overfitting with limited support triplets. When fixing $M=32$ and varying $N\in\{3,5,10\}$, we observe that selecting slightly more experts (e.g., 5 vs. 3) leads to improved performance, yet increasing $N$ to 10 causes only a slight drop, suggesting that MoEMeta is relatively insensitive to $N$ within a reasonable range.